OpenStudy (anonymous):

A=2pi*r^2+2pi*rh, solve for positive r Please unlock my intuitions after I divided both sides of the equation by 2pi, and got this: hr+r^2=A/(2pi) Thank you.

OpenStudy (anonymous):

A=2pi*r^2+2pi*rh, hr+r^2=A/(2pi) now use the quadratic formula here r^2 +hr-A/(2pi)=0..a=1,b=h,c=A/2pi using only positive value r=[-h+sqrt(h^2 -(4A/2pi))]/2 answer......